Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards)
A. rolling a sum of 4 on two fair six sided dice
B. drawing a black seven out of a standard 52 card deck given it's not a face card of an Ace.
C. rolling a sum of 11 on two fair six sided dice.
D. rolling a sum of 7 on two fair six sided dice.
E. drawing a three out of a standard 52 card deck given it's not a face card or an Ace.

Answer by ewatrrr(24785)

 
Hi,
Note:
B and E refer to standard 52 card deck/12 face cards and 4 aces. (52-16)= 36
Probability of:
A. rolling a sum of 4 on two fair six sided dice = 3/36 = 1/12
B. drawing a black seven,given it's not a face card of an Ace = 2/36 = 1/18
C. rolling a sum of 11 on two fair six sided dice = 2/36 = 1/18
D. rolling a sum of 7 on two fair six sided dice = 6/36 = 1/6
E. drawing a three, given it's not a face card or an Ace = 4/36 = 1/9
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12